OpenStudy (liziekay):

Add as indicated. (52° 45') + (36° 55')

OpenStudy (liziekay):

I got 89 degrees, but I can't figure the minutes out..

zepdrix (zepdrix):

Adding degrees with degrees, minutes with minutes,\[\large\rm 88^o100'\]That seems like a sensible first step, ya? :)

OpenStudy (liziekay):

yes!

zepdrix (zepdrix):

60 minutes is 1 degree. So we can borrow 60 of those minutes to increase our degrees to 89.

OpenStudy (liziekay):

i think I got it, would it be 40 minutes?

zepdrix (zepdrix):

\(\large\rm 89^o40'\) 40 left over from the 100? Yayy good job \c:/

OpenStudy (liziekay):

Convert the following to degrees and minutes. 19.15° what about this one, I know it's 19 degrees

zepdrix (zepdrix):

Well we know that 1 degree is 60 minutes. So if we wanted to convert something like say 4 degrees to minutes, we would multiply the 4 by 60, yes?\[\large\rm 4^o\quad=\quad (4\cdot 60)'\quad=\quad240'\]

OpenStudy (liziekay):

yes

zepdrix (zepdrix):

So that's what we want to do with our decimal remainder. We'll take this fraction of a degree, and multiply it by 60, our conversion value. \(\large\rm 19.15^o\quad=\quad 19^o(.15\cdot60)'\)

OpenStudy (liziekay):

so basically, everytime you take the remainder and multiply by 60?

OpenStudy (liziekay):

and that would be 9

zepdrix (zepdrix):

\[\large\rm 19^o9'\]For that type of problem, yes. And keep in mind, the decimal value is `less than 1 full degree`, so you should always end up with `less than 60 minutes` when multiplying. That's a good way to check your work I suppose. If you have something 24.89 and somehow that .89 turns into 85 minutes, then you know you've made a mistake because it's larger than 60.

OpenStudy (liziekay):

I totally get it now, thanks!